OpenStudy (haleyelizabeth2017):

Find the polar equation in rectangular form. r=-16sin theta Please no direct answers....I just need the formula(s) for doing this (:

myininaya (myininaya):

Hint multiply both sides by r.

myininaya (myininaya):

And then try to use some or all of the following: \[r^2=x^2+y^2 \\ r \cos(\theta)=x \\ r \sin(\theta)=y \\ \tan(\theta)=\frac{y}{x}\]

OpenStudy (haleyelizabeth2017):

@myininaya...I'm not sure what it is supposed to look like when multiplying r by (-16 sin theta)....

OpenStudy (haleyelizabeth2017):

Do you have an example by chance doing this?

myininaya (myininaya):

just r*(-16sin(theta)) or -16 r sin(theta) since multiplication is associative and commutative and don't forgot you also need to multiply the other side by r

OpenStudy (haleyelizabeth2017):

Okay, So it would be \[r^2=-16r~sin\theta\]

myininaya (myininaya):

yes

myininaya (myininaya):

now use the equations mentioned by me or the ones in the picture provided by @Dani_Rose (they are the same )

OpenStudy (haleyelizabeth2017):

Will I need to isolate the r then? rather than it being r^2?

myininaya (myininaya):

you are trying to write in terms of x and y... r^2 is...

myininaya (myininaya):

r sin(theta) is ...

OpenStudy (haleyelizabeth2017):

r sin theta would then be equal to the square root of (x^2+y^2), right?

OpenStudy (haleyelizabeth2017):

wait...I think I'm going at that wrong.

myininaya (myininaya):

did you read the equations I posted

OpenStudy (haleyelizabeth2017):

oh...sorry r sin theta is y.

myininaya (myininaya):

these: \[r^2=x^2+y^2 \\ r \cos(\theta)=x \\ r \sin(\theta)=y \\ \tan(\theta)=\frac{y}{x} \] you are using those to write \[r^2=-16 r \sin(\theta) \text{ in terms of } x \text{ and } y\]

myininaya (myininaya):

you see that you can replace r^2 (the thing on the left hand side) with? you see the r sin(theta) on the right hand side which can be replaced with?

myininaya (myininaya):

and yes r sin(theta) is y ...so just replace that with y and replace r^2 with...

OpenStudy (haleyelizabeth2017):

okay, r^2 would be sqrt(x^2+y^2)

myininaya (myininaya):

do you see the first equation I wrote?

myininaya (myininaya):

r^2=x^2+y^2 ...

myininaya (myininaya):

you shouldn't be involving a square root in that

OpenStudy (haleyelizabeth2017):

So sorry, I don't know where my brain is today ;-;

myininaya (myininaya):

so replace the r^2 and r sin(theta) with x^2+y^2 and y respectively and you are done writing as a Cartesian equation

OpenStudy (haleyelizabeth2017):

so \[x^2+y^2=-16y\]?

myininaya (myininaya):

yep

OpenStudy (haleyelizabeth2017):

ah..

myininaya (myininaya):

any questions?

OpenStudy (haleyelizabeth2017):

Well, I was wondering how to get it into one of these forms...one second

myininaya (myininaya):

this is a circle you can write in the form (x-h)^2+(y-k)^2=r^2

myininaya (myininaya):

involves completing square

OpenStudy (haleyelizabeth2017):

Ohhh, maybe that's what it is

myininaya (myininaya):

\[x^2+y^2+16y=0\] we don't have to worry with the x-stuff because the x part as already been written as a square we need to do something about the y part of this...

OpenStudy (haleyelizabeth2017):

OH!

OpenStudy (haleyelizabeth2017):

I love this stuff...

OpenStudy (haleyelizabeth2017):

so we just complete the square of \[y^2+16y+~~=0\]

myininaya (myininaya):

what do we need to add to both sides to complete the square there

OpenStudy (haleyelizabeth2017):

so (y+8)^2=64 right?

myininaya (myininaya):

right so with out equation we need to retrice64 on both sides

OpenStudy (haleyelizabeth2017):

Because you divide by 2 and then square that...add to both sides

myininaya (myininaya):

\[x^2+y^2+16y=0 \\ x^2+y^2+16y+64=64 \\ x^2+(y+8)^2=64\]

OpenStudy (haleyelizabeth2017):

Awesome, that makes a ton more sense. Thank you!!!

myininaya (myininaya):

np